# Enhancing Study-Level Inference from Clinical Trial Papers via RL-based Numeric Reasoning

### v1

@@ Abstract
Enhancing Study-Level Inference from Clinical Trial Papers via RL-based Numeric Reasoning
@@ 1 Introduction
Systematic reviews are the cornerstone of evidence-based medicine, offering rigorous syntheses of available studies to guide clinical decision-making Murad et al. (2016). A critical component of systematic reviews is the extraction of study-level numeric evidence (e.g., event counts or standard deviations) from one or multiple corresponding clinical trial papers and derive the conclusions for each outcome and comparison under assessment. However, automating this extraction remains an open challenge. As illustrated in Figure 1 (prior work), previous studies Pronesti et al. (2025) has primarily framed this task as a retrieval-based question answering problem: given a query about an outcome, systems retrieve relevant study fragments and infer conclusions based on the retrieved text.
@@ 3.2.1 SFT with CoT
We first adapt a pretrained language model to our task via supervised fine-tuning (SFT), which adapts a pretrained language model $\pi_{\theta}$ to reflect a domain-specific distribution $\mathcal{P}$. This is achieved by minimising the negative log-likelihood over a dataset of example sequences, encouraging the model to increase the probability of desired outputs. In our context, the goal is to enable the model to map free-text descriptions of study results to structured outputs that capture outcome information in a standardised schema. Each training example consists of a biomedical study, paired with the corresponding outcome of interest (Table 6). The output is represented in YAML format and encodes either binary outcomes (with event counts and totals for intervention and comparator groups) or continuous outcomes (with group-wise means, standard deviations, and sample sizes). The model learns to generate these structured summaries conditioned on the corresponding textual evidence. That is, given prompt–target pairs $(\mathbf{x},\mathbf{y})$, the loss is computed as:
@@ 3.2.2 RL with Fine-grained Rewards
In our setup, the model acts as a policy $\pi_{\theta}$ that takes as input a textual passage describing clinical study results and outputs a structured response. Each response consists of a thought process enclosed in a <think> tag, followed by a YAML object encoding outcome data, using the same schema adopted in SFT. For each passage $x$, multiple candidate completions $\{y_{i}\}_{i=1}^{G}$ are sampled from the model to encourage robustness and diversity. These completions are scored using rule-based reward functions that evaluate factual correctness and adherence to format. The raw rewards $R_{i}$ are then normalised across the group:
$A_{i}=\frac{R_{i}-\mathbb{E}[R_{j}]}{\sqrt{\mathbb{V}[R_{j}]}}$
The policy is optimised using a clipped, KL-regularised objective that encourages agreement with high-reward behaviours while maintaining proximity to a reference model $\pi_{\mathrm{ref}}$:
$\mathcal{L}_{\text{GRPO}}(\theta)=\;$ $\mathbb{E}\bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\min\Big{(}rA_{i},\,\text{clip}(r,1{-}\epsilon,1{+}\epsilon)A_{i}\Big{)}$
$\quad-\beta\,\text{KL}[\pi_{\theta}\,\|\,\pi_{\text{ref}}]\bigg{]}$
where $r=p_{i,t}(\theta)$ is the token-level probability ratio, and $\beta$ governs the regularisation strength.
@@
where $v_{j}\approx\hat{v}_{j}$ means exact match for integers and absolute difference $<10^{-3}$ for floats. If parsing fails or types mismatch, we set $R_{\mathrm{CR}}=0$.
@@
This reward ensures that the output adheres to a valid YAML schema.
@@
Final Reward. The final reward is a weighted combination of format and correctness components:
@@ 4.1 Experimental setup
EIR=#errors that flip the conclusion#extraction $\text{EIR}=\frac{\#\text{errors that flip the conclusion}}{\#\text{extraction errors}}$
@@ 4.2 Main Results
Fine-tuning improves performance significantly. Qwen2.5-7B-SFT outperforms all pretrained models of similar or larger size. The Qwen2.5-7B-RL model establishes a new state-of-the-art on both datasets, surpassing all baselines—including the 405B model—in nearly every metric on RCTs and closely matching it on CochraneForest, despite being nearly 58x smaller. In addition, when compared directly to its pretrained counterpart, Qwen2.5-7B-RL shows large gains in EM (+21.2), accuracy (+16.1), and F1 (+16.7) on the RCTs dataset, highlighting the effectiveness of RL.
@@ 4.4 Analyses
We present a qualitative example from CochraneForest (Figure 8, in appendix), comparing the reasoning and outputs of Qwen2.5-7B-SFT and Qwen2.5-7B-RL on a challenging case where the exact values for the comparator and outcome are not explicitly reported in the biomedical study. While the SFT model fails to extract the relevant information, the RL model demonstrates stronger reasoning capabilities: it identifies population percentages from one of the study’s tables and correctly maps them to the total number of participants previously extracted, thereby inferring the desired values. Notably, even our human annotators initially struggled to locate the correct numbers in this scenario.
@@ 5 Related work
Lehman et al. (2019) demonstrated that accurately extracting evidence is the primary bottleneck for determining treatment efficacy. O’Doherty et al. (2024) showed using abstracts alone rather than structured PICO data can worsen synthesis quality, underscoring the importance of data representation. Our model processes the full content of the paper, providing a more comprehensive basis.
@@
While LLMs have shown promise in biomedical text mining, they face challenges synthesising complex evidence and numerical reasoning. Shaib et al. (2023) found that GPT-3 struggled with multi-document synthesis and biased effect reporting. Nye et al. (2020) introduced a framework to map evidence and infer conclusions. Yun et al. (2024) assessed how effectively LLMs can extract numerical data, noting challenges with continuous outcomes and distinguishing similar measures.
@@ 6 Conclusion
In this paper we presented a quantitative reasoning framework for automating evidence extraction in systematic reviews, shifting away from shallow textual inference toward structured numeric understanding. By directly modelling the process experts follow, we move closer to automating a key step in evidence-based medicine. Our proposed system, combining SFT and RL numeric extraction models with reasoning over effect estimates, significantly outperforms retrieval-based baselines and even large-scale language models on both the CochraneForest and the RCTs benchmarks. These results affirm the value of reasoning-driven supervision for complex scientific tasks and point to RL as a promising avenue for developing reliable and domain-aligned extraction systems in evidence-based medicine.
@@ 7 Limitations
While this work makes meaningful contributions toward reasoning-based evidence extraction, several limitations remain. First, the system assumes that all relevant numerical evidence is explicitly reported and cleanly extractable, which is often not the case in real-world clinical studies. Missing values, indirect reporting, and inconsistent formats can hinder reliable extraction. At present, the model does not attempt to identify or flag missing information, which limits its utility in incomplete or noisy settings. Training models to detect and explicitly report missing or uncertain values is a critical direction for future development. Additionally, our approach currently focuses on a limited set of outcome types and uses predefined logical rules, which may constrain generalisation to more complex or diverse clinical scenarios. Addressing these limitations will be key to deploying robust, scalable systems for real-world systematic review automation.
@@ Appendix B Prompts
The prompts used for synthetic data annotation and for training are shown in Figure 5 and 6, respectively. For training, a temperature of 0.7 and 2,048 tokens as maximum output length are used.
@@ Appendix E Hyperparameters and APIs
The weights of the final reward model have been set to maximise the importance of the correctness reward (0.8) while still ensuring the model follows the YAML format reward (0.1) and the already learnt thought format reward present in the base model (0.1). A tuning of such weights revealed that assigning equal weights to the 3 components leads to a degradation of the performances.

### v2

# Enhancing Study-Level Inference from Clinical Trial Papers via Reinforcement Learning-Based Numeric Reasoning

@@ Abstract
Enhancing Study-Level Inference from Clinical Trial Papers via Reinforcement Learning-Based Numeric Reasoning
@@ 1 Introduction
Systematic reviews are the cornerstone of evidence-based medicine, offering rigorous syntheses of available studies to guide clinical decision-making Murad et al. (2016). A critical component of systematic reviews is the extraction of study-level numeric evidence (e.g., event counts or standard deviations) from one or multiple corresponding clinical trial papers and derive the conclusions for each outcome and comparison under assessment. However, automating this extraction remains an open challenge. As illustrated in Figure 1 (prior work), previous studies Pronesti et al. (2025) have primarily framed this task as a retrieval-based question answering problem: given a query about an outcome, systems retrieve relevant study fragments and infer conclusions based on the retrieved text.
@@
0252550507575100100525254545656585860606262646466666868Retrieval Precision (%)F1 ScoreGPT-4 (0125)GPT-35-turbo (0125)LLaMa 3.3 70BQwen 2.5 72B
@@ 3.2.1 SFT with CoT
We first adapt a pretrained language model to our task via supervised fine-tuning (SFT), which adapts a pretrained language model $\pi_{\theta}$ to reflect a domain-specific distribution $\mathcal{P}$. This is achieved by minimising the negative log-likelihood over a dataset of example sequences, encouraging the model to increase the probability of desired outputs. In our context, the goal is to enable the model to map free-text descriptions of study results to structured outputs that capture outcome information in a standardised schema. Each training example consists of a biomedical study, paired with the corresponding outcome of interest (Table 6). The output is represented in YAML format and encodes either binary outcomes (with event counts and totals for intervention and comparator groups) or continuous outcomes (with group-wise means, standard deviations, and sample sizes). The model learns to generate these structured summaries conditioned on the corresponding textual evidence. That is, given prompt-target pairs $(\mathbf{x},\mathbf{y})\sim\mathcal{P}$, the loss is computed as:
@@ 3.2.2 RL with Fine-grained Rewards
In our setup, the model acts as a policy $\pi_{\theta}$ that takes as input a textual passage describing clinical study results and outputs a structured response. Each response consists of a thought process enclosed in a <think> tag, followed by a YAML object encoding outcome data, using the same schema adopted in SFT. For each passage $\mathbf{x}$, $G$ candidate completions $\{y_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot\mid\mathbf{x})$ are sampled from the reference policy $\pi_{\text{old}}$ to encourage robustness and diversity. These completions are scored using rule-based reward functions that evaluate factual correctness and adherence to format. The raw rewards $R_{i}$ are then normalised across the group:
$A_{i}=\frac{R_{i}-\mathbb{E}[R_{j}]}{\sqrt{\mathbb{V}[R_{j}]}},\quad j\in\{i,...,G\}$
where $\mathbb{E}[R_{j}]$ and $\mathbb{V}[R_{j}]$ are respectively the mean and variance of the rewards for the group of responses. The policy is optimised using a clipped, KL-regularised objective that encourages agreement with high-reward behaviours while maintaining proximity to a reference model $\pi_{\mathrm{ref}}$:
$\mathcal{L}_{\text{GRPO}}(\theta)=\;$ $\mathbb{E}\bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Big(p_{i,t}(\theta)A_{i},\,$
$\quad\hskip-28.45274pt\text{clip}(p_{i,t}(\theta),1-\varepsilon,1+\varepsilon)A_{i}\Big)-\beta\,\text{KL}[\pi_{\theta}\,\|\,\pi_{\text{ref}}]\bigg]$
where $\beta$ governs the regularisation strength and $p_{i,t}(\theta)$ is the token-level probability ratio defined as follows:
$p_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid x,y_{i,<t})}$
@@
where $v_{j}\approx\hat{v}_{j}$ means exact match for integers and absolute difference $10^{-3}$ for floats. If parsing fails or types mismatch, we set $R_{\mathrm{CR}}=0$.
@@
This reward ensures that the output adheres to a valid YAML schema for the predicted outcome type.
@@
Final Reward.
The final reward is a weighted combination of format and correctness components:
@@ 4.1 Experimental setup
$\text{EIR}=\frac{\#\text{errors that flip the conclusion}}{\#\text{extraction errors}}$
@@ 4.2 Main Results
Fine-tuning improves performance significantly. Qwen2.5-7B-SFT outperforms all pretrained models of similar or larger size. The Qwen2.5-7B-RL model establishes a new state-of-the-art on both datasets, surpassing all baselines—including the 405B model—in nearly every metric on RCTs and closely matching it on CochraneForest, despite being nearly 58x smaller. In addition, when compared directly to its pretrained counterpart, Qwen2.5-7B-RL shows large gains in EM (+20.9), accuracy (+16.1), and F1 (+16.7) on the RCTs dataset, highlighting the effectiveness of RL.
@@ 4.4 Analyses
Reward Dynamics.
Figure 3 shows the reward dynamics during RL training. Thought Format Reward (TFR) and Format Reward (FR) increase rapidly and plateau early, indicating that the model quickly learns the reasoning scaffold and YAML schema. This behaviour also suggests that the model inherently learns to predict the correct outcome type at an early stage. By contrast, Correctness Reward (CR) improves more gradually and continues rising after TFR/FR have stabilised, showing that numerical correctness requires longer training. These dynamics are consistent with our main results and design choices: structural aspects are mastered quickly, whereas correctness is more demanding and depends on the model’s ability to apply reasoning across each component of the final output. Overall, the reward dynamics illustrate how reinforcement signals guide the model from surface-level alignment toward substantive reasoning ability.
01001002002003003004004005005006006007007000.20.20.40.40.60.60.80.811StepTraining RewardYAML Format RewardThought Format RewardCorrectness Reward
Figure 3: Rewards dynamics. Format rewards plateau quickly during training, whereas the correctness reward improves more gradually, suggesting that LLMs naturally learn to follow structural patterns while still improving on quality rewards.
@@
We present a qualitative example from CochraneForest (Figure 9, in appendix), comparing the reasoning and outputs of Qwen2.5-7B-SFT and Qwen2.5-7B-RL on a challenging case where the exact values for the comparator and outcome are not explicitly reported in the biomedical study. While the SFT model fails to extract the relevant information, the RL model demonstrates stronger reasoning capabilities: it identifies population percentages from one of the study’s tables and correctly maps them to the total number of participants previously extracted, thereby analytically inferring the desired values. Notably, even our human annotators initially struggled to locate the correct numbers in this scenario.
@@ 5 Related work
Lehman et al. (2019) demonstrated that accurately extracting evidence is the primary bottleneck for determining treatment efficacy. O’Doherty et al. (2024) showed using abstracts alone rather than structured PICO data can worsen synthesis quality, underscoring the importance of data representation. Our model processes the full content of the paper, providing a more comprehensive basis. Our work can be seen as an example of scientific argument mining at the global discourse level Al Khatib et al. (2021) and contributes to the broader agenda of AI for Science Eger et al. (2025).
@@
While LLMs have shown promise in biomedical text mining, they face challenges synthesising complex evidence and numerical reasoning. Shaib et al. (2023) found that GPT-3 struggled with multi-document synthesis and biased effect reporting. Nye et al. (2020) introduced a framework to map evidence and infer conclusions. Yun et al. (2024) assessed how effectively LLMs can extract numerical data, noting challenges with continuous outcomes and distinguishing similar measures. Wang et al. (2025) introduced a method that integrates LLMs with code generation to extract clinical study outcomes. However, their evaluation was restricted to selected cancer-related reviews, and the reliance on structured prompts and domain-specific heuristics raises questions about scalability to broader therapeutic areas. Lai et al. (2025) evaluated LLMs for data extraction and risk of bias assessment in complementary medicine, reporting high performance, but their focus on a relatively narrow and domain-specific set of trials limits the generalizability of the findings. Similarly, Sun et al. (2024) assessed LLMs for automated data extraction from randomized trials, finding encouraging results, yet performance was uneven across outcome types, with continuous measures proving particularly challenging. These studies highlight that while LLMs show high potential, existing work focuses on narrow domains, curated benchmarks, or simplified tasks, leaving open the question of how well they generalise to the diverse evidence encountered in real-world systematic reviews.
Reinforcement Learning for Numerical Evidence Extraction.
Previous work has proven that SFT can be used to improve models such as BERT Devlin et al. (2019) for biomedical text mining Lee et al. (2019); Xie et al. (2022). Reinforcement learning has improved model alignment for complex reasoning objectives (Wei et al., 2022; Lambert et al., 2024; Guo et al., 2025). Recent work Lin et al. (2025) demonstrates that RL with verifiable rewards can improve model performance on clinical reasoning tasks like EHR-based calculations and trial matching. To the best of our knowledge, we are the first to design and apply RL-based approaches specifically to the extraction of numerical evidence from biomedical studies for use in systematic reviews.
@@ 6 Conclusion
In this paper we presented a quantitative reasoning framework for automating evidence extraction in systematic reviews, shifting away from shallow textual inference toward structured numeric understanding. By directly modelling the process domain experts follow, we move closer to automating a key step in evidence-based medicine. Our proposed system, combining SFT and RL numeric extraction models with reasoning over effect estimates, significantly outperforms retrieval-based baselines and even large-scale language models on both the CochraneForest and the RCTs benchmarks. These results affirm the value of reasoning-driven supervision for complex scientific tasks and point to RL as a promising avenue for developing reliable and domain-aligned extraction systems in evidence-based medicine.
Beyond improving performance, our approach highlights the importance of aligning machine reasoning with expert workflows, ensuring interpretability and trustworthiness. Looking forward, this work opens opportunities for integrating automated evidence extraction into clinical decision support pipelines, ultimately reducing the manual burden on researchers and accelerating the translation of medical knowledge into practice. Future research may extend this framework to other scientific domains where quantitative reasoning is core.
@@ 7 Limitations
While this work makes meaningful contributions toward reasoning-based evidence extraction, several limitations remain. First, the system assumes that all relevant numerical evidence is explicitly or implicitly reported and cleanly extractable, which is often not the case in real-world clinical studies. Missing values and inconsistent formats can hinder reliable extraction. At present, the model does not attempt to identify or flag missing information, which limits its utility in incomplete or noisy settings. Training models to detect and explicitly report missing or uncertain values is a critical direction for future development. Additionally, our approach currently focuses on a limited set of outcome types and uses predefined logical rules, which may constrain generalisation to more complex or diverse clinical scenarios. Addressing these limitations will be key to deploying robust systems for real-world systematic review automation.
@@ Appendix B Prompts
The prompts used for synthetic data annotation and for training are shown in Figure 6 and 7, respectively. For training, a temperature of 0.7 and 2,048 tokens as maximum output length are used.
@@ Appendix E Hyperparameters and APIs
The weights of the final reward model have been set to maximise the importance of the correctness reward (0.8) while still ensuring the model follows the YAML format reward (0.1) and the already learnt thought format reward present in the base model (0.1). A tuning of such weights revealed that assigning equal weights to the 3 components leads to a degradation of the performance.